OpenStudy (anonymous):

The length of an altitude of an equilateral triangle is 12 feet. Find the length of a side of the triangle.

OpenStudy (anonymous):

\[ H = \frac {\sqrt{3}} {2} a \]

OpenStudy (anonymous):

side length \( = 8 \sqrt{3} \)